Minutes PSB Upgrade WG Meeting 200, 12th of September 2017
Participants: S. Albright, J-B. Bonnamy, J.D. Devine, G.P. Di Giovanni, E. Fortescue-Beck, R. Froeschl, G.M. Georgiev, G. Guidoboni, K. Hanke, J. Hansen, C. Hessler, B. Mikulec, A. Newborough, S. Pittet, F. Schmidt, J. Tan, W. Weterings.
Approval of Minutes
The minutes of the LIU-PSB WG #199 were approved.
Communications
- LIU Project Team meeting (24th of August 2017
).
- Chamonix meeting. It will take place from the 29th of January to the 2nd of February and be focused on LIU and HL-LHC projects.
- Storage and space management until the end of LS2. The requested space is much less than what was initially foreseen. Have all the requests been submitted? A carefull check is needed, including the start and end date of the requests.
- Report on Linac4 H- source exchange from Alessandra Lombardi
Schedule of LIU and Equipment groups dedicated meetings is available here
.
- LIU and EN/HE joint meeting (5th of September
).
- PSBooster tools are ALL ready and confirmed today by J-B. Bonnamy.
- Requests for handling the scrapers have been presented.
- To be followed up. Currently there are 2 months of delay in the budget.
- LIU and Be-BI meeting (21st of September
).
- J. Tan presented the activity of the Booster. Under control
- 2 requests for BPMs (1 in BI and 1 in BTP line).
- Good budget consumption.
- ECRs to be produced for the new PUs (dirven by ABT).
- Vacuum Test for the Head and Tail Dumps (1st of September 2017).
- Procedures and responsibilities were clarified.
Follow-up of Open Actions
- PSB integration (J.M. Lacroix, B. Riffaud): C. Hessler corrected the discrepancies found with intergration and J.M. Lacroix has now the updated information. Work in progress.
- M. Calviani/I. Lamas: "Prepare the ECR for the PSB scraper/absorber describing both the dismantling of the current equipment and the installation of the new one." Deadline to be postponed to the next meeting .
- S. Pittet: "Provide document with the final specifications for ALG1+2 power converter upgrade." The paper is curculating. The action can be closed.
- K. Hanke: "Discuss with SMB the consolidation of building 141." . K. Hanke could not access the buildig but he managed to take few pictures available here
. He cannot judge if consolidation is required: G. Petrika should follow this up and if necessary organize a visit. Action Closed for K. Hanke.
- C. Hessler: "Send the new MAD-X file to I. Efthymiopoulos and T. Birtwistle for the upgrade of the LIU-PSB and current MAD-X optics, and also the layout database." Postponed by 1 month (middle of October).
- A. Newborough: "Provide the final price of the recombination dipoles and correctors." More news about prices will be available by the end of the month.
A reminder should be sent to:
- T. Dobers: "Make a list of starting points of the transfer lines to be sent for approval in an EDMS document." (action expired on the 31st of August).
- G. Petrika:
- "Prepare a document summarizing the ventilation needs for LIU-PSB." (action expired on the 31st of August).
- New: "Verify the consolidation required for building 141." Due date 26th of September.
- "Iterate with integration on the installation of the new piping system for cooling." (due date 30th of September)
- "Release a new version of the EDMS document 1758628, summarizing the cooling needs for LIU-PSB.2 (due date 30th of September)
- "Check the viability of the removal of the obsolete heater units during the YETS during LS2." (due date 30th of September) .
- L.A.Lopez-Hernandez: "Report about the status of the building 245 modifications." (due date 30th of September).
AOB
Follow up on additional BPMs in the PSB Transfer Lines request
- 1 BPM in the BI line + BPM in the BTP line (see talk
).
- LIU approval obtained. ABT will drive the activity in collaboration with OP, BI and ACE and report back to LIU-PSB
- ABT is responsible for this activity, including ECRs, DICs, etc... .
- The work units have been created: J. Tan is the responsible and S. Prodon will move the money there.
Comments:
- B. Mikulec: 1 coordinator from ABT is needed. Proposal for the ECRs:
- An ECR for the injection line exists already. A new ECR for BPM in BI should be prepared.
- An ECR for the extraction lines does not exist yet. A chapter concerning the new BPM should be included.
Tour de table
- S. Pittet would like to include a short review on the injection Q-strips at at the next LIU-PSB working group meeting. An AOB of 5-10 minutes would be sufficient.
- E. Fortescue-Beck reminded that the deadline for YETS hardware requests is END of September.
- B. Mikulec mentioned an issue with the injection watchdog. 16 OASIS signals were requested but the current hardware has not a large enough capability. An additional installation might be necessary.
- J. Hansen has contacted S. Prodon to discuss the budget but not answer received yet. He asked to A. Newborough if he can start the integration of the vacuum chambers: A. Newborough confirmed that few new magnet modules have been provided (at their best knowledge). Drawings were already approved and available.
- G.M. Georgiev asked for a meeting with CO and B. Mikulec in order to the define the installation priorities for the YETS. G.M. Georgiev will need this information before the 28th of September.
- F. Schmidt is working on the digitization of the "old" PSB and PS documentation and he found a paper from the 1968 about the dump resistors. To be checked if the numbers there are still valid. Work in progress.
- B. Mikulec asked to A. Newborough if he knows the measurments that M. Buzio needs to do, but A. Newborough has not heard anything about it.
Next LIU-PSB meeting: 26th of September.
There are minutes attached to this event.
Show them.